A commuter train blows its 200 Hz horn as it approaches a crossing. The speed of sound is 330 m/s.

(a) An observer waiting at the crossing receives a frequency of 210 Hz. What is the speed of the train (in m/s)?

(b) What frequency (in Hz) does the observer receive as the train moves away?
Answer #1

a) fobs / fs = [vw / vw - vs]

210 / 200 = 330 / (330 - vs )

1.05 = 330 / (330 - vs )

(330 - vs ) = 330 / 1.05

vs = 15.7 m/s

b) fobs = f0 [vw / vw + vs]

= 200 [330 / 330 + 15.7]

= 191 Hz